punisher432:

The formula for the circumference of a circle is 4072-01-01-03-00_files/i0150000.jpg. Determine the circumference when the radius r is 10 cm.

dude:

Just substitute r into the equation given \(C=2\pi r\) \(C=2\pi (10)\)

punisher432:

so 20 cm?

punisher432:

or 20piecm

dude:

20 pi

punisher432:

why is it the pi one if i may ask?

dude:

\(C=2π(10)\) The original equation has a pi 2*10 is 20 (You have to keep the pi) \(20\pi\) or 62.8315...

punisher432:

oh gotcha thank you bud

dude:

np lol